OpenStudy (anonymous):

integral sqrt(x^2+2x+2)/x

OpenStudy (anonymous):

what it's supposed to look like (for clarification) http://www.wolframalpha.com/input/?i=integral+sqrt%28x^2%2B2x%2B2%29%2Fx

OpenStudy (anonymous):

i've tried everything i could think of for values to substitute

OpenStudy (amistre64):

yeah, nice arcsinh in there :) try completing the square under the radical? then you can attempt a trig sub perhaps

OpenStudy (anonymous):

i did taht so i got sqrt((x+1)^2+1)/x but from there what do i do? what do i trig sub for?

OpenStudy (amistre64):

well, off hand i know that tan^2+1 = sec^2

OpenStudy (anonymous):

right

OpenStudy (amistre64):

lets (x+1) = tan(u)

OpenStudy (anonymous):

i tried that and then i ended up getting the integral of (secx)^3 / tanx-1 http://www.wolframalpha.com/input/?i=integral+sec^3x%2F%28tanx-1%29

OpenStudy (anonymous):

sorry, in this case x should be u

OpenStudy (anonymous):

and then from there idk what to do

OpenStudy (amistre64):

\[\int \sec(u)\frac{\sec^2(u)}{\tan(u)+1}du\]that at least has ln integral sitting in it; my thoughth is a by part perhaps?

OpenStudy (anonymous):

"ln integral" what do you mean by that? can you point that out for me

OpenStudy (amistre64):

sec^2 is the derivative of tan+1: u'/u integrates to ln u

OpenStudy (anonymous):

so do you think i should do integration by parts?

OpenStudy (amistre64):

well, its just sitting there ... might as well give it a shot :)

OpenStudy (anonymous):

hm i still dont know how to make use of it

OpenStudy (amistre64):

\[\sec(u)~\ln(tan(u)+1)-\int \ln(\tan(u)+1)~\sec(u)\tan(u)~du\]

OpenStudy (anonymous):

did you achieve this result via integration by parts?

OpenStudy (anonymous):

if so, what were your terms?

OpenStudy (amistre64):

well, since the integration was already set; i used sec^2/(tan+1) as dv, and sec as u

OpenStudy (amistre64):

not sure of that int v du is workable tho ....

OpenStudy (anonymous):

hm ok i understand the integration by parts. let me try to work out the second integral now haha

OpenStudy (amistre64):

.... im pretty sure its messier than the first :)

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

well through a bunch of other various sdbustitutions

OpenStudy (anonymous):

i arrived at the integral of sqrt(u^2+1)/(u-1) http://www.wolframalpha.com/input/?i=integral+of+sqrt%28u^2%2B1%29%2F%28u-1%29

OpenStudy (amistre64):

i wonder if there some multiplier of 1 that would make that sqrt vanish

OpenStudy (anonymous):

if we can solve that simpler one then we solve the original question

OpenStudy (anonymous):

try working on the new one to see. i got stuck on that one also

OpenStudy (amistre64):

\[\frac{\sqrt{x^2+2x+2}}{x}\] \[\frac{x^2+2x+2}{x\sqrt{x^2+2x+2}}\] \[\frac{x+2}{\sqrt{x^2+2x+2}}+\frac{2}{x\sqrt{x^2+2x+2}}\] \[\frac22\frac{x+2}{\sqrt{x^2+2x+2}}+\frac{2}{x\sqrt{x^2+2x+2}}\]

OpenStudy (amistre64):

the left part is simple enough, thats another ln up; the right part is looking triggy now

OpenStudy (anonymous):

WOAH

OpenStudy (amistre64):

lol, not ln but sqrt

OpenStudy (anonymous):

ok so the integral of 1/sqrt(x^2+1) is that sinhx?

OpenStudy (anonymous):

i think so, right?

OpenStudy (amistre64):

we never really covered the hyperbolics in my class so im not as familiar with them

OpenStudy (anonymous):

arsinh***

OpenStudy (amistre64):

\[2\sqrt{x^2+2x+2}~+~2\int\frac{1}{x\sqrt{(x+1)^2+1}}dx\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

if we let u = 1/x and dv = 1/sqrt((x+1)^2+1) dx

OpenStudy (anonymous):

u = -1/x^2 and v = inverse sinh(x+1)

OpenStudy (anonymous):

du = -1/x^2 dx*

OpenStudy (anonymous):

hm well i guess i'll close this question for now. i think simplifying it to that last integral was a lot more help than i was expecting. thank you so much amistre64 :)

OpenStudy (amistre64):

youre welcome :)